A 0.02 kg rock strikes the ground with 0.36 J of kinetic energy. What was its velocity?


1
Expert's answer
2021-02-05T01:31:57-0500
"KE=\dfrac{1}{2}mv^2,""v=\sqrt{\dfrac{2KE}{m}}=\sqrt{\dfrac{2\cdot0.36\ J}{0.02\ kg}}=6\ \dfrac{m}{s}."